# Temporally Consistent Object-Centric Learning by Contrasting Slots

### v1

@@ 5 Conclusion
Limitations of our work include the fixed number of slots during initialization. Additionally, we cannot directly control the segmentation granularity of entities. For further limitations and failure cases are discussed in App. K.
@@ Appendix C Implementation of Slot-Slot Contrastive Loss
Appendix D Dataset Details
@@
Appendix E Metrics Details
@@
To differentiate between per-frame (image-based) and video-wide evaluations, we use ”Image” as a prefix for the metrics (e.g., Image FG-ARI and Image mBO) when computed on individual frames. When we do not use an additional prefix, we refer to the ”Video” version of the same metric when computed across entire videos. We are particularly interested in video-based metrics, as they additionally consider the consistency of object masks.
Appendix F Baseline Details
@@
For the YouTube-VIS 2021 dataset, the authors of VideoSAUR provided results for both configurations, so we directly used the available checkpoints. However, for the MOVi datasets, results and model for VideoSAUR trained with DINOv2 features were not available. Therefore, we trained VideoSAUR with the default configuration using DINOv2 features.
@@
Appendix G Per-frame Scene Decomposition
@@
Appendix H Instance-Awareness of Dense Features
@@ Appendix I SlotFormer
Appendix J Details and Visual Examples on MOVI-C Occluded
@@
Appendix K Limitations and Failure Cases
@@
Appendix L Additional Examples

### v2

@@ 5 Conclusion
Limitations of our work include the fixed number of slots during initialization. Additionally, we cannot directly control the segmentation granularity of entities. Further limitations and failure cases are discussed in App. L.
@@ Appendix C Implementation of Slot-Slot Contrastive Loss
Appendix D Feature Reconstruction Loss as Regularizer
To promote better object discovery we also use feature reconstruction loss. Feature reconstruction loss, $\mathcal{L}_{\text{rec}}$, measures the discrepancy between the predicted features $\hat{h}_{t}$ and the true features $h_{t}$ at each time step $t$. In our case the features correspond to self-supervised DINOv2 features. The loss could be computed using a common distance metric such as Mean Squared Error (MSE):
$\mathcal{L}_{\text{rec}}=\sum_{t=1}^{T-1}||h_{t}-\hat{h}_{t}||^{2}$ (S2)
The loss also serves as an effective regularizer, mitigating undesired behaviors that can arise from the contrastive nature of slot-slot contrastive loss. For example, slot-slot contrastiveloss can’t pull slots representing different objects together because it is minimized alongside the feature reconstruction loss $\mathcal{L}_{\text{rec}}$. This way, we maximize slot-slot similarity while still requiring each slot to be informative about original inputs. So region-wise reconstruction with an MLP decoder decoding slots individually is an effective regularizer, preventing “wrong slots pulling” behavior as otherwise pulled slots will not contain the information about the object they are responsible to reconstruct.
Another key scenario is when an object disappears. In this case, it is important to understand what happens to the corresponding slot and how its behavior is governed by the objectives. In that case, we want the corresponding slot to maintain object information. Given the additional reconstruction loss, it is possible by ignoring the disappeared object’s slot (thus serving as latent memory until object reappearance). This behavior is evident in the Fig. 6 showing fewer active slots compared to baseline that uses all the available slots.
Appendix E Dataset Details
@@
Appendix F Metrics Details
@@
To differentiate between per-frame (image-based) and video-wide evaluations, we use "Image" as a prefix for the metrics (e.g., Image FG-ARI and Image mBO) when computed on individual frames. When we do not use an additional prefix, we refer to the "Video" version of the same metric when computed across entire videos. We are particularly interested in video-based metrics, as they additionally consider the consistency of object masks.
Appendix G Baseline Details
VideoSAUR
@@ Appendix G Baseline Details
For the YouTube-VIS 2021 dataset, the authors of VideoSAUR provided results for both configurations, so we directly used the available checkpoints. However, for the MOVi datasets, results and model for VideoSAUR trained with DINOv2 features were not available. Therefore, we trained VideoSAUR with the default configuration( matching the resolution with Slot Contrast) using DINOv2 features.
@@
SAM2
To compare how close current object-centric methods are to supervised methods we compared Slot Contrast with SAM2 as a supervised zero-shot baseline for temporal consistency. As SAM2 is trained on a large dataset with dense video annotations ($190.9K$ masklets), using its tracking can improve segmentation consistency (limited to objects discovered in the first frame). However, while SAM2 can be used only for object tracking, our method is not limited to tracking; it jointly does both object discovery in videos and learns consistent object representations with their masks. We evaluate SAM2’s tracking capabilities by combining SAM2 with initial frame object discovery using video-based DINOSAUR (i.e, feature reconstruction objective on videos) and Slot Contrast object discovery (see Table S2). We show that Slot Contrast halves the gap between unsupervised object-centric learning and zero-shot SAM2 ($5.5$ vs $12.3$ FG-ARI), while using Slot Contrast object discovery is helpful for overall tracking with SAM2 ($+2.8$ FG-ARI).
Table S2: Temporal consistency on YouTube-VIS 2021.
Feat. Rec. + SAM2 Slot Contrast + SAM2 VideoSAURv2 Slot Contrast FG-ARI 43.5 46.3 31.2 38.0 mBO 40.9 43.7 29.7 33.7
Figure S2: SlotContrast vs SAM2 tracking. SAM2 is limited to track only objects that appeared and discovered in the first frame.
In addition, in Fig. S2, we show limitation of such baseline: detecting and tracking later appearing objects due to missing initial masks. Evaluating SAM2 on YTVIS’s first-frame objects gives $46.3$ mBO ($+6\%$), while for the later-appearing objects, mBO drops to $7.82$ ($-34.48\%$).This highlights SAM2’s strength in tracking first-frame objects and its limitation in detecting and tracking later objects due to missing initial masks.
SAVI++
We compared Slot Contrast with weakly supervised method SAVi++. We used improved SAVi similar to VideoSAUR (see App. C.5 VideoSAUR), reaching $42.8$ FG-ARI on MOVi-E. In contrast, unconditioned optical-flow SAVi and depth SAVi++ are only $28.1$ and $31.7$ as reported by Bao et al. [3]. While adding depth signal in SAVi++ could be treated as weak supervision, it indeed improves SAVi $16.0$ mBO, reaching $22.1$ mBO, but still lagging behind both VideoSAUR and SlotContrast.
Appendix H Per-frame Scene Decomposition
@@
Appendix I Instance-Awareness of Dense Features
@@
Appendix K Details and Visual Examples on MOVI-C Occluded
@@
Appendix L Limitations and Failure Cases
@@
Appendix M Additional Examples